OpenStudy (anonymous):

any simplification for: (0.67*0.67*0.67-0.001)/(0.67*0.67+0.067+0.01)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

let a=.67, b=.1 so u have (a^3-b^3)/(a^2+ab+b^2)=a-b=.57